OpenStudy (anonymous):

3. Mr. Richards needs to price items that must be sold within 90 days, and he has two options. In Option A, Mr. Richards prices items at $225 and gives $1.00 off the price for every day that the item does not sell. In Option B, he prices items at $250 and gives $2.00 off for every day the item does not sell. Write an equation for each option that relates the price of the item to the number of days that the item has not sold.

OpenStudy (anonymous):

ah!

OpenStudy (anonymous):

225-1.00?

hartnn (hartnn):

225p-i*1=90......i think

OpenStudy (aripotta):

225 - x = p where x is the amount of days not sold and p is the price...i think that would work for the first one

hartnn (hartnn):

is that an option?

OpenStudy (aripotta):

and then 250 - 2x = p

OpenStudy (anonymous):

no i did that as my own answer last night but i deleted it i dont think it was right

OpenStudy (aripotta):

who, me?

OpenStudy (anonymous):

no mine lol

OpenStudy (anonymous):

the 225p i

hartnn (hartnn):

is it 225 for each item or total?

hartnn (hartnn):

if total, then @AriPotta is correct

hartnn (hartnn):

if its for each then u need to have 225p

OpenStudy (anonymous):

thanks Aripotta and Hartnn